OpenStudy (anonymous):

A loan of $27000 is made at 3.75% interest, compounded annually. After how many years will the amount due reach $43,000 or more? please help! thank you!

OpenStudy (anonymous):

M = P( 1 + i )^n 43000 = 27000(1 + .0375)^n

OpenStudy (anonymous):

solve for n gives ...

OpenStudy (anonymous):

about 12.64

OpenStudy (anonymous):

A principal of $1200 is invested at 5.75% interest, compounded annually. How many years will it take to accumulate $2000 or more to the account. help :(

OpenStudy (anonymous):

2k in total gain or over the 1.2 original?

OpenStudy (anonymous):

2000 or more in the account. 1200 is invested

OpenStudy (anonymous):

its the same interest formula I would think 2000 = 1200(1 + .0575)^n

OpenStudy (anonymous):

n is about 9.14